# ReQISC: A Reconfigurable Quantum Computer Microarchitecture and Compiler Co-Design

### v1

@@ Abstract.
Despite remarkable milestones in quantum computing, the performance of current quantum hardware remains limited. One critical path to higher performance is to expand the quantum ISA with basis gates that have higher fidelity and greater synthesis capabilities than the standard $\mathrm{CNOT}$. However, this substantially increases gate calibration overhead and introduces challenges in compiler optimization. Consequently, although more expressive ISAs (even complex, continuous gate sets) have been proposed, they are still primarily proof-of-concept and have not been widely adopted.
To move beyond these hurdles and unlock the performance gains offered by expressive continuous ISAs, we introduce the concept of “reconfigurable quantum instruction set computers” (ReQISC), which incorporates: (1) a unified microarchitecture capable of directly implementing arbitrary 2Q gates equivalently, i.e., $\mathrm{SU}(4)$ modulo 1Q rotations, with theoretically optimal gate durations given any 2Q coupling Hamiltonians; (2) a compilation framework tailored to ReQISC primitives for end-to-end synthesis and optimization, comprising a program-aware pass that refines high-level representations, a program-agnostic pass for aggressive circuit-level optimization, and an $\mathrm{SU}(4)$-aware routing pass that minimizes hardware mapping overhead.
We detail the hardware implementation to demonstrate the feasibility, in terms of both pulse control and calibration of this superior gate scheme on realistic hardware. By leveraging the expressivity of $\mathrm{SU}(4)$ and the time minimality realized by the underlying microarchitecture, the $\mathrm{SU}(4)$-based ISA achieves remarkable performance, with a $4.97$-fold reduction in average pulse duration to implement arbitrary 2Q gates, compared to the usual $\mathrm{CNOT}/\mathrm{CZ}$ scheme on mainstream flux-tunable transmons. Supported by the end-to-end compiler, ReQISC outperforms the conventional $\mathrm{CNOT}$-ISA, SOTA compiler, and pulse implementation counterparts, in significantly reducing 2Q gate counts, circuit depth, pulse duration, qubit mapping overhead, and program fidelity losses. For the first time, ReQISC makes the theoretical benefits of continuous ISAs practically feasible.
* Corresponding author: chenjianxin@tsinghua.edu.cn.
@@ 1. Introduction
Quantum hardware has advanced rapidly over the past two decades, achieving milestones including quantum supremacy (arute2019quantum), quantum utility (kim2023evidence), and going below the fault-tolerant threshold (acharya2024quantum). Today’s quantum processors come with detailed device specifications such as native gate sets and pulse-level controls, accessible through a cloud interface or a complete system delivery. However, hardware noise presents critical challenges for both noisy intermediate-scale quantum (NISQ) (preskill2018quantum) and fault-tolerant application-scale quantum (FASQ) (preskill2025beyond) devices, significantly limiting executable program scales and impacting computational reliability (arute2019quantum; kim2023evidence; acharya2024quantum).
Figure 1. Advancing architectural support for quantum advantage.
@@
To pursue a practical quantum advantage, many sophisticated architectural supports have been proposed by exploring advanced instruction set architectures (ISA) beyond the conventional $\mathrm{CNOT}$-based paradigm (Figure 1). For specific applications, novel 2Q quantum gates like Google’s Sycamore from the $\mathrm{fSim}$ family offer higher fidelity and difficulty to simulate classically (arute2019quantum), while gates like $\mathrm{SQiSW}$ aim to improve fidelity and synthesis efficiency (huang2023quantum). Physical platforms developed by Google, Rigetti, IonQ, Quantinuum, and IBM have also demonstrated continuous gate sets (foxen2020demonstrating; abrams2020implementation; ionqPartialGates; ibmFractionalGates; quantinuumArbitraryAngleGates), which appear to yield improvements in fidelity (wright2019benchmarking; nam2020ground). At the frontier, recent works have proposed using the entire $\mathbf{SU}(4)$ group as the most expressive ISA, achieving optimal-duration gates on flux-tunable transmons with $99.37\%$ average fidelity in experiments (chen2024one; chen2025efficient).
@@
Adopting more expressive quantum ISAs offers significant theoretical benefits, most notably a reduction in the 2Q gate count required for circuit synthesis (abrams2020implementation; foxen2020demonstrating; chen2024one; ionqPartialGates; ibmFractionalGates). This is demonstrated by comparing the Haar-random synthesis cost of representative 2Q basis gates against the standard $\mathrm{CNOT}$-equivalent gates. For instance, the cost is $3$ for $\mathrm{CNOT}/\mathrm{CZ}/\mathrm{iSWAP}$ (shende2004minimal), around $2.21$ for $\mathrm{SQiSW}$ (huang2023quantum), $2$ for $\mathrm{B}$ gate (zhang2004minimum), and even lower for combinatorial or continuous gate sets (peterson2020fixed). However, despite the variety of quantum ISAs available, none have successfully usurped the dominance of the $\mathrm{CNOT}$-based ISA due to several inherent challenges:
@@
Hardware implementation and calibration. Implementing versatile 2Q gates with high precision and simple control via unified gate schemes across diverse quantum platforms remains a significant challenge, especially for complex, continuous quantum ISAs. For example, although $\mathrm{B}$ gate was proposed decades ago (zhang2004minimum), it was only recently natively implemented (chen2025efficient; nguyen2024programmable); IBM’s fractional $\mathrm{ZZ}$ gates are deployed on their Heron QPUs (ibmFractionalGates), but they have the same gate time as $\mathrm{CZ}$ and on average $1.5$ times the error rate of $\mathrm{CZ}$, according to public data from IBM Quantum Platform (ibmFractionalGates). Furthermore, each 2Q gate must be carefully and periodically calibrated to ensure high precision. While proposals exist to mitigate calibration overhead (kelly2014optimal), their effectiveness is not guaranteed. Thus, a program with numerous distinct 2Q gates is hard to execute in practice.
@@
Compilation strategy. Although the capability to implement more complex ISAs has been demonstrated, effectively utilizing them to compile programs into fewer gates with reduced circuit depth remains largely unaddressed. Neither gate set transpilation (peterson2022optimal; mckinney2024mirage; yale2024noise) nor numerically optimal synthesis (lao2021designing; kalloor2024quantum) effectively exploits their synthesis potential for real-world program compilation. We observe that ISA-customized strategies are crucial at every compilation stage, from high-level synthesis to hardware mapping. Otherwise, a theoretically superior ISA may be inferior in practice (kalloor2024quantum).
@@
To overcome these challenges and establish the practical superiority of expressive continuous ISAs, we introduce ReQISC (Reconfigurable Quantum Instruction Set Computers). ReQISC is a full-stack framework that integrates compilation, ISA design, and pulse control to boost performance across versatile quantum hardwares. At its core, ReQISC utilizes the most expressive $\mathbf{SU}(4)$ as its ISA (i.e., all 2Q gates) and makes it the most performant with minimal gate time through our proposed microarchitecture (gate scheme). The ReQISC microarchitecture can straightforwardly implement arbitrary 2Q gates up to local equivalence in optimal time, via simple pulse controls, on any 2Q coupling Hamiltonian. Recent experimental progress has affirmed the viability of specific mechanisms in this direction (chen2025efficient); our scheme, however, represents a more generalized approach with extensive architectural exploration, extending the advantages to a broader spectrum of hardware. Our key contributions are as follows:
❶We introduce a unified microarchitecture for native, time-optimal realization for all of $\mathbf{SU}(4)$ with arbitrary coupling Hamiltonians. We validate the pulse-control simplicity and calibration feasibility on representative hardwares.
❷We overcome the difficulty of implementing near-identity 2Q gates (chen2024one), by replacing them with their mirrors (cross2019validating) and subsequently altering qubit mapping at compile time, ensuring no extra 2Q gate count overhead.
@@
❸We present the first end-to-end compiler tailored to $\mathrm{SU}(4)$ characteristics. Based on the lower-bound analysis of required $\mathrm{SU}(4)$ resource for synthesizing multi-qubit circuits, our compiler employs both program-aware and program-agnostic passes to optimize the overall $\mathrm{SU}(4)$ gate count. We address the program-specific calibration overhead and also effectively manage the trade-off between calibration overhead and aggressive 2Q gate count reduction.
@@
❹We propose the program-aware template-based synthesis method that first pre-synthesizes optimal $\mathrm{SU}(4)$-based circuit templates for refined intermediate representations (IR) and then selectively assembles the whole circuit.
@@
❺We also propose a hierarchical, program-agnostic synthesis pass to aggressively lower the 2Q gate count, building on circuit partitioning and approximate synthesis. We introduce the partitioning compactness metric to guide co-optimization and develop a directed acyclic graph (DAG) compacting pass based on approximate commutation rules in our findings.
@@
❻Furthermore, we introduce a novel $\mathrm{SU}(4)$-aware routing pass, dubbed mirroring-SABRE, which significantly reduces the routing overhead engendered by topology constraints.
@@ 2.1. Quantum instruction set
Similar to its classical counterpart, a quantum ISA serves as an interface between software and hardware, by mapping high-level semantics of quantum programs to low-level native quantum operations or pulse sequences on hardware. Quantum instructions involve physical operations on qubits that alter their states. Typically, these instructions include qubit initialization, a universal gate set, and measurement. The universal gate set is the key component of a quantum ISA that dominates its hardware-implementation accuracy and cost, as well as the software expressivity. It is well known that a gate set consisting of any 2Q entangling gate (e.g., $\mathrm{CNOT}$, $\mathrm{iSWAP}$) together with all possible 1Q gates can achieve universal computation (bremner2002practical). Instructions from a universal gate set correspond to dynamics governed by the system Hamiltonian $H$, potentially influenced by external drives. The evolution of a quantum state over time $t$ is described by the unitary operator $U(t)=e^{-iHt}$ which realizes the desired quantum gate up to a global phase.
@@ 2.2. Canonical decomposition and the Weyl chamber
The group $\mathbf{SU}(2^{n})$ is a real manifold with dimension $4^{n}-1$, each of which corresponds to an $n$-qubit unitary. A generic 2Q gate within $\mathbf{SU}(4)$, despite having $15$ real parameters, can have its nonlocal behavior fully characterized by only $3$ real parameters. This method, known as Canonical decomposition or KAK decomposition from Lie algebra theory, is widely adopted in quantum computing (zhang2003geometric; tucci2005introduction; bullock2003arbitrary; zulehner2019compiling).
For any $U\in\mathbf{SU}(4)$, there exists a unique vector $\vec{\eta}=(x,y,z)\in W\subseteq\mathbb{R}^{3}$, along with $V_{1},V_{2},V_{3},V_{4}\in\mathbf{SU}(2)$ and a global phase, such that
$U=g\cdot(V_{1}\otimes V_{2})e^{-i\vec{\eta}\cdot\vec{\Sigma}}(V_{3}\otimes V_{4}),\,g\in\{1,i\}$
where $\vec{\Sigma}\equiv(XX,YY,ZZ)$ (tucci2005introduction). The set
@@
is known as the Weyl chamber (zhang2003geometric), and $\vec{\eta}\in W$ is known as the Weyl coordinate of $U$. We also refer to a gate of the form $\mathrm{Can}(x,y,z)\coloneqq e^{-i\vec{\eta}\cdot\vec{\Sigma}}$ as a canonical gate. Two 2Q gates $U$ and $V$ are considered locally equivalent if they differ only by 1Q gates. For example, $\mathrm{CNOT}\sim\mathrm{CZ}$, as $\mathrm{CZ}=(I\otimes H)\,\mathrm{CNOT}\,(I\otimes H)$ and their canonical coordinates are both $(\pi/4,0,0)$.
@@ 3. Our Proposal: ReQISC
Figure 2. ReQISC workflow. It incorporates (1) the most performant (time-optimal realization) microarchitecture to engineer any arbitrary $\mathrm{SU}(4)$ gates given the hardware coupling Hamiltonian and (2) an end-to-end compilation framework with three-stage optimization passes to generate executable $\mathrm{SU}(4)$-based circuits. The $\mathrm{SU}(4)$ ISA is expressed as the $\{\mathrm{Can},\,\mathrm{U}3\}$ gate set.
Figure 3. The ReQISC microarchitecture (gate scheme) directly implements arbitrary 2Q gates in a unified control. Common 2Q coupling form and standard 1Q Rabi drives with control parameters $\Omega_{1}$, $\Omega_{2}$, $\delta$, $\tau$, are sufficient to generate an evolution locally equivalent to the target unitary $U$. With the additional 1Q corrections appended/prepended to the solved evolution $e^{-i\tau(H+H_{1}+H_{2})}$, the circuit in the right figure exactly implements $U$.

### v2

# Reconfigurable Quantum Instruction Set Computers for High Performance Attainable on Hardware

@@ Abstract.
Despite remarkable milestones in quantum computing, the performance of current quantum hardware remains limited. One critical path to higher performance is to expand the quantum ISA with basis gates that have higher fidelity and greater synthesis capabilities than the standard $\mathrm{CNOT}$. However, this substantially increases gate calibration overhead and introduces challenges in compiler optimization. Consequently, although more expressive ISAs (even complex, continuous gate sets) have been proposed, they still remain primarily proofs-of-concept and have not been widely adopted.
To move beyond these hurdles and unlock the performance gains offered by expressive continuous ISAs, we introduce the concept of “reconfigurable quantum instruction set computers” (ReQISC). It incorporates (1) a unified microarchitecture capable of directly implementing arbitrary 2Q gates equivalently, i.e., $\mathrm{SU}(4)$ modulo 1Q rotations, with theoretically optimal gate durations given any 2Q coupling Hamiltonian and (2) a compilation framework tailored to ReQISC primitives for end-to-end synthesis and optimization, comprising a program-aware pass that refines high-level representations, a program-agnostic pass for aggressive circuit-level optimization, and an $\mathrm{SU}(4)$-aware routing pass that minimizes hardware mapping overhead.
We detail the hardware implementation to demonstrate the feasibility of this superior gate scheme in terms of both pulse control and calibration. By leveraging the expressivity of $\mathrm{SU}(4)$ and the time minimality realized by the underlying microarchitecture, the $\mathrm{SU}(4)$-based ISA achieves remarkable performance, with a $4.97$-fold reduction in average pulse duration to implement arbitrary 2Q gates, compared to the usual $\mathrm{CNOT}/\mathrm{CZ}$ scheme on mainstream flux-tunable transmons. Supported by the end-to-end compiler, ReQISC outperforms the conventional $\mathrm{CNOT}$-based ISA, state-of-the-art compiler, and pulse implementation counterparts by significantly reducing 2Q gate count, circuit depth, pulse duration, qubit mapping overhead, and program fidelity losses. For the first time, ReQISC makes the theoretical benefits of continuous ISAs practically feasible.
@@ 1. Introduction
Figure 1. Advancing architectural support for practical quantum advantage.
@@
Quantum hardware has advanced rapidly over the past two decades, achieving remarkable milestones including quantum supremacy (Arute et al., 2019), quantum utility (Kim et al., 2023), and achieving error rates below the fault-tolerant threshold (Acharya et al., 2024). Today’s quantum processors come with detailed device specifications such as native gate sets and pulse-level controls, accessible through a cloud interface or a complete system delivery. However, hardware noise presents critical challenges for both noisy intermediate-scale quantum (NISQ) (Preskill, 2018) and fault-tolerant application-scale quantum (FASQ) (Preskill, 2025) devices, significantly limiting executable program scales and impacting computational reliability (Arute et al., 2019; Kim et al., 2023; Acharya et al., 2024).
@@
To pursue a practical quantum advantage, many sophisticated architectural supports have been proposed by exploring advanced instruction set architectures (ISA) beyond the conventional $\mathrm{CNOT}$-based paradigm (Figure 1). For specific applications, novel 2Q quantum gates like Google’s Sycamore from the $\mathrm{fSim}$ family offer higher fidelity and are harder to simulate classically (Arute et al., 2019), while gates like $\mathrm{SQiSW}$ aim to improve fidelity and synthesis efficiency (Huang et al., 2023). Physical platforms developed by Google, Rigetti, IonQ, Quantinuum, and IBM have also demonstrated continuous gate sets (Foxen et al., 2020; Abrams et al., 2020; IonQ, 2023; IBM Quantum, 2024; Quantinuum, 2024), which appear to yield improvements in fidelity (Wright et al., 2019; Nam et al., 2020). At the frontier, recent works have proposed using the entire $\mathbf{SU}(4)$ group as the most expressive ISA, achieving optimal-duration gates on flux-tunable transmons with $99.37\%$ average fidelity in experiments (Chen et al., 2024, 2025).
@@
Adopting more expressive quantum ISAs offers significant theoretical benefits, most notably a reduction in the 2Q gate count required for circuit synthesis (Abrams et al., 2020; Foxen et al., 2020; Chen et al., 2024; IonQ, 2023; IBM Quantum, 2024). This is demonstrated by comparing the Haar-random synthesis cost of representative 2Q basis gates against the standard $\mathrm{CNOT}$-equivalent gates. For instance, the cost is $3$ for $\mathrm{CNOT}/\mathrm{CZ}/\mathrm{iSWAP}$ (Shende et al., 2004), around $2.21$ for $\mathrm{SQiSW}$ (Huang et al., 2023), $2$ for $\mathrm{B}$ gate (Zhang et al., 2004), and even lower for combinatorial or continuous gate sets (Peterson et al., 2020). However, despite the variety of quantum ISAs available, none have successfully challenged the dominance of the $\mathrm{CNOT}$-based ISA due to several inherent challenges:
@@
Hardware implementation and calibration. Implementing diverse 2Q gates with high precision and simple control via unified gate schemes across diverse quantum platforms remains a significant challenge, especially for complex, continuous quantum ISAs. For example, although $\mathrm{B}$ gate was proposed decades ago (Zhang et al., 2004), it has only recently been implemented natively (Chen et al., 2025; Nguyen et al., 2024); IBM’s fractional $\mathrm{ZZ}$ gates are deployed on their Heron QPUs (IBM Quantum, 2024), but they have the same gate time as $\mathrm{CZ}$ and on average $1.5$ times the error rate of $\mathrm{CZ}$, according to public data from IBM Quantum Platform (IBM Quantum, 2024). Furthermore, each 2Q gate must be carefully and periodically calibrated to ensure high precision. While proposals exist to mitigate calibration overhead (Kelly et al., 2014), their effectiveness is not guaranteed. Thus, a program with numerous distinct 2Q gates presents significant execution challenges in practice.
@@
Compilation strategy. Although the capability to implement more complex ISAs has been demonstrated, effectively utilizing them to compile programs into fewer gates with reduced circuit depth remains largely unaddressed. Neither gate set transpilation (Peterson et al., 2022; McKinney et al., 2024; Yale et al., 2024) nor numerically optimal synthesis (Lao et al., 2021; Kalloor et al., 2024) effectively exploits their synthesis potential for real-world program compilation. We observe that ISA-customized strategies are crucial at every compilation stage, from high-level synthesis to hardware mapping. Otherwise, a theoretically superior ISA may be inferior in practice (Kalloor et al., 2024).
@@
To overcome these challenges and establish the practical superiority of expressive continuous ISAs, we introduce ReQISC (Reconfigurable Quantum Instruction Set Computers). ReQISC is a full-stack framework that integrates compilation, ISA design, and pulse control to boost performance across various quantum hardware platforms. At its core, ReQISC utilizes the most expressive $\mathbf{SU}(4)$ as its ISA (i.e., all 2Q gates) and makes it the highest-performing with minimal gate time through our proposed microarchitecture (gate scheme). The ReQISC microarchitecture can straightforwardly implement arbitrary 2Q gates up to local equivalence in optimal time, via simple pulse controls, given any 2Q coupling Hamiltonian. Recent experimental progress has affirmed the viability of specific mechanisms in this direction (Chen et al., 2025); our scheme, however, represents a more generalized approach with extensive architectural exploration, extending the advantages to a broader hardware spectrum. Our key contributions are:
❶We introduce a unified microarchitecture for the native, time-optimal realization of the full $\mathrm{SU}(4)$ group under arbitrary coupling Hamiltonians. We validate the pulse-control simplicity and calibration feasibility on representative hardware platforms. To ensure experimental feasibility across all 2Q gates, we incorporate a gate mirroring mechanism that resolves the control singularity of near-identity gates at compile time without incurring 2Q gate count overhead.
@@
❷We present the first end-to-end compiler tailored to $\mathrm{SU}(4)$ characteristics. Based on the lower-bound analysis of required $\mathrm{SU}(4)$ resources for synthesizing multi-qubit circuits, our compiler employs both program-aware and program-agnostic passes to optimize the overall $\mathrm{SU}(4)$ gate count. We address the program-specific calibration overhead and also effectively manage the trade-off between calibration overhead and aggressive 2Q gate count reduction.
@@
❸We propose the program-aware template-based synthesis method that first pre-synthesizes optimal $\mathrm{SU}(4)$-based circuit templates for refined intermediate representations (IR) and then selectively assembles the whole circuit.
@@
❹We also propose a hierarchical, program-agnostic synthesis pass to aggressively lower the 2Q gate count, building on circuit partitioning and approximate synthesis. We introduce the partitioning compactness metric to guide co-optimization and develop a directed acyclic graph (DAG) compacting pass based on approximate commutation rules.
@@
❺Furthermore, we introduce a novel $\mathrm{SU}(4)$-aware routing pass, dubbed mirroring-SABRE, which significantly reduces the routing overhead induced by topology constraints.
@@ 2.1. Quantum instruction set
Similar to its classical counterpart, a quantum ISA serves as an interface between software and hardware, by mapping the high-level semantics of quantum programs to low-level native quantum operations or pulse sequences on hardware. Quantum instructions involve physical operations on qubits that alter their states. Typically, these instructions include qubit initialization, a universal gate set, and measurement. The universal gate set is the key component of a quantum ISA that dominates its hardware-implementation accuracy and cost, as well as software expressivity. It is well known that a gate set consisting of any 2Q entangling gate (e.g., $\mathrm{CNOT}$, $\mathrm{iSWAP}$) together with all possible 1Q gates can achieve universal computation (Bremner et al., 2002). Instructions from a universal gate set correspond to dynamics governed by the system Hamiltonian $H$, potentially influenced by external drives. The evolution of a quantum state over time $t$ is described by the unitary operator $U(t)=e^{-iHt}$ which realizes the desired quantum gate up to a global phase.
@@ 2.2. Canonical decomposition and the Weyl chamber
The group $\mathbf{SU}(2^{n})$ is a real manifold with dimension $4^{n}-1$; each point in this manifold corresponds to an $n$-qubit unitary. A generic 2Q gate within $\mathbf{SU}(4)$, despite having $15$ real parameters, can have its nonlocal behavior fully characterized by only $3$ real parameters. This method, known as Canonical decomposition or KAK decomposition from Lie algebra theory, is widely adopted in quantum computing (Zhang et al., 2003; Tucci, 2005; Bullock and Markov, 2003; Zulehner and Wille, 2019).
Any $U\in\mathbf{SU}(4)$ can be defined by a unique vector $\vec{\eta}=(x,y,z)\in W\subseteq\mathbb{R}^{3}$, along with $V_{1},V_{2},V_{3},V_{4}\in\mathbf{SU}(2)$, s.t.,
(1) $U=g\cdot(V_{1}\otimes V_{2})e^{-i\vec{\eta}\cdot\vec{\Sigma}}(V_{3}\otimes V_{4}),\,g\in\{1,i\}$
where $\vec{\Sigma}\equiv(XX,YY,ZZ)$ (Tucci, 2005). The set
@@
is known as the Weyl chamber (Zhang et al., 2003), and $\vec{\eta}\in W$ is known as the Weyl coordinate of $U$. We also refer to a gate of the form $\mathrm{Can}(x,y,z)\coloneqq e^{-i\vec{\eta}\cdot\vec{\Sigma}}$ as a canonical gate. Two 2Q gates $U$ and $V$ are considered locally equivalent if they differ only by 1Q gates, and we write $U\sim V$ to denote the equivalence relation. For example, $\mathrm{CNOT}\sim\mathrm{CZ}$, as $\mathrm{CZ}=(I\otimes H)\,\mathrm{CNOT}\,(I\otimes H)$ and their canonical coordinates are both $(\pi/4,0,0)$.
Figure 2. ReQISC workflow. It incorporates (1) the most performant (time-optimal realization) microarchitecture to engineer any arbitrary $\mathrm{SU}(4)$ given the hardware coupling Hamiltonian and (2) an end-to-end compilation framework with three-stage optimization passes to generate executable $\mathrm{SU}(4)$-based circuits. The $\mathrm{SU}(4)$ ISA is expressed as the $\{\mathrm{Can},\,\mathrm{U}3\}$ gate set.
@@ 3. Our Proposal: ReQISC
(a) Landscape of ReQISC microarchitecture (gate scheme)
(b) ReQISC’s micro-control logic from the system perspective
Figure 3. ReQISC microarchitecture (gate scheme) directly implements arbitrary 2Q gates with a unified control. Common 2Q coupling Hamiltonians and standard 1Q Rabi drives with control parameters $\Omega_{1}$, $\Omega_{2}$, $\delta$, $\tau$, are sufficient to generate an evolution locally equivalent to the target unitary $U$.